OpenStudy (anonymous):

Evaluate. 2⁶/4³ A.) 16/3 B.) 3/16 C.) 1/2 D.) 1

OpenStudy (anonymous):



OpenStudy (anonymous):

Thanks (:

OpenStudy (anonymous):

4^3 = 2^2^3 = 2^6, so 2^6/2^6 = 1